OpenStudy (anonymous):

1. Factor: x2 + 11x + 10 (x + 10)(x +1) (x + 11)(x + 10) (x + 2)(x + 5) prime

OpenStudy (across):

What is so hard about these?

OpenStudy (bahrom7893):

10 and 1

OpenStudy (bahrom7893):

factors of c that add up to b right?

OpenStudy (hoblos):

x2 + 11x + 10 = x2 +x+10x+10 = x(x+1) +10(x+1) = (X+1)(X+10)